# New Algebraic Points on Curves

### v1

@@ Abstract.
Let $C/\mathbb{Q}$ be a smooth projective absolutely irreducible curve of genus $\geq 2$. For an extension $L/\mathbb{Q}$, let $C(L)_{\mathrm{new}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{P\in C(L)\;:\;\mathbb{Q}(P)=L\}$. Let $n\geq 2$. We conjecture that $C(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of degree $n$ number fields $L$ when ordered by absolute discriminant. We prove our conjecture for $n=3$ with $C$ any of the modular curves $X_{0}(23)$, $X_{0}(29)$, $X_{0}(31)$ and $X_{0}(64)$. Moreover, we prove the analogue of this conjecture for the unit equation, again with $n=3$.
@@ 1. Introduction
Let $C$ be a smooth projective and absolutely irreducible curve of genus $g$, defined over a number field $K$. Let $D$ be a reduced effective divisor on $C$ and consider the punctured curve X=CDX=C-D. The Euler characteristic of $X$ is given by $\chi(X)=2-2g-\deg(D)$. We say that $X$ is hyperbolic if $\chi(X)<0$. We let $\mathscr{X}/\mathcal{O}_{K}$ be a model for $X$. The famous Faltings–Siegel Theorem asserts that $\mathscr{X}(\mathcal{O}_{K})$ is finite. We mention three special cases.
@@
  • Let $X=\mathbb{P}^{1}-\{0,1,\infty\}$, which has Euler characteristic 1-1 and is therefore hyperbolic. It can be shown that $\varepsilon\in\mathscr{X}(\mathcal{O}_{K})$ if and only if $\varepsilon$ and $1-\varepsilon$ are units. Thus Faltings–Siegel generalises Siegel’s Theorem on the unit equation, i.e., the unit equation
@@
A typical instance of the above question is when $X=C$ is a complete curve of degree $g\geq 2$, and $\mathcal{L}$ is the set of all extensions $L/K$ of a fixed degree $n$. This instance of the question is well-considered in literature; for an excellent survey we recommend [25], but merely mention two famous results.
@@
  • A theorem of Harris and Silverman [12] asserts that a (complete) curve $C/\mathbb{Q}$ of genus $\geq 2$ has infinitely many quadratic points if and only if $C$ is hyperelliptic or bielliptic. This theorem builds on a deep theorem of Faltings asserting that if $A/K$ is an abelian variety and $V$ is a subvariety not containing any abelian variety then $V(K)$ is finite.
In this paper we offer a conjectural statistical (partial) response to Question 1. Let $X/K$ be a hyperbolic curve as before, and let $L/K$ be an extension. The conjecture is concerned which what we call the set of $L$-new points that appear over $L$ on $X/K$; we define these by
@@
We note that the $L$-new points are ones defined over $L$, but not over any strictly smaller extension. Before stating our conjecture, we explain what we mean by a family $\mathcal{L}$ of extensions $L/K$.
@@
Let $K$ be a number field. Let $n\geq 2$. Let $G_{1},\dotsc,G_{r}$ be transitive subgroups of $S_{n}$. Let $\upsilon_{1},\dotsc,\upsilon_{s}$ be a finite number of places of $K$, and let $L_{i}/K_{\upsilon_{i}}$ be degree $n$ étale algebras. By the degree $n$ family $\mathcal{L}$ with data $(G_{1},\dotsc,G_{r};L_{1},\dotsc,L_{s})$ we mean all extensions $L/K$ such that the following hold.
@@
If $X=C$ is a smooth projective absolutely irreducible curve over a number field $K$, then
@@
  • It is tempting to conjecture that $\mathscr{X}(\mathcal{O}_{L})=\mathscr{X}(\mathcal{O}_{K})$ for $100\%$ of $L\in\mathcal{L}$. This however is false. For example, let $\mathcal{L}$ be all quartic number fields $L/\mathbb{Q}$. Let $X/\mathbb{Q}$ be a hyperbolic curve such that $\mathscr{X}(\mathbb{Z})=\emptyset$, but $\mathscr{X}(\mathbb{Z}[(1+\sqrt{5})/2])\neq\emptyset$. A positive proportion of quartic number fields $L$, when ordered by absolute discriminant, contain $\mathbb{Q}(\sqrt{5})$ ([4, Corollary 1.2]) and therefore $\mathscr{X}(\mathcal{O}_{L})\neq\mathscr{X}(\mathbb{Z})$ for those fields.
@@
The purpose of this paper is to prove Conjecture 2 for several hyperbolic curves, with $\mathcal{L}$ the set of all cubic fields. All these hyperbolic curves have infinitely many cubic points.
@@
To motivate what comes next, we mention the following “single source theorem” due to the authors [14, Theorem 1.1]. In this context, by a primitive degree $n$ point on $C/K$, we mean an algebraic point $P\in C(\overline{K})$ such that $K(P)/K$ has degree $n$, and the Galois group of its Galois closure is a primitive subgroup of $S_{n}$.
Theorem.
This theorem is called the “Single Source Theorem” since, with finitely many exceptions, all primitive degree $n$ points come from a single source which is the morphism $\varphi:C\rightarrow\mathbb{P}^{1}$. In view of our Conjecture 2, it makes sense to consider the family of degree $n$ number fields we obtain from the fibres of this morphism, and to attempt to show that this family contains $0\%$ of degree $n$ number fields. We shall give a result in this spirit.
Let $C$ be a smooth projective absolutely irreducible curve over $K$, and let $\varphi:C\rightarrow\mathbb{P}^{1}$ be a morphism of degree $n\geq 2$. Then the curve $C$ possesses a plane affine (possibly singular) model
C:F(T,Y)=0C\;:\;F(T,Y)=0
where $F(T,Y)\in\mathbb{Q}(T)[Y]$ has degree $n$ in $Y$, is irreducible over $\overline{\mathbb{Q}}(T)$, and where the morphism $\varphi$ is given by the map $(T,Y)\mapsto T$. Hilbert’s irreducibility theorem asserts the existence of a thin set $\Theta\subset\mathbb{Q}$, such that for $t\in\mathbb{Q}-\Theta$, the polynomial $F(t,Y)\in\mathbb{Q}[Y]$ is irreducible of degree $n$. Thus $C$ has infinitely many degree $n$ points. For $t\in\mathbb{Q}-\Theta$ we write $K_{t}$ for the degree $n$ number field defined by $F(t,Y)=0$, and let $\Delta(\mathcal{O}_{K_{t}})$ be the discriminant of the ring of integers $\mathcal{O}_{K_{t}}$ of $K_{t}$. We define the discriminant set of $F$ to be
$\operatorname{Disc}_{F}\;=\;\{\lvert\Delta(\mathcal{O}_{K_{t}})\rvert\;:\;t\in\mathbb{Q}-\Theta\}.$
@@
where $g\in\mathbb{Q}[T]$ is a squarefree polynomial and $h(T)\in\mathbb{Q}(T)$ is a rational function. Suppose $g$ has even degree. Let $G=\operatorname{Gal}(g)$. Suppose that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$. Then there is some $0<\alpha<1$ such that, as $X\rightarrow\infty$,
@@ 2. Counting integers with certain restricted factorisations
Let $\mathbb{P}$ be the set of prime numbers and let $\mathcal{P}\subseteq\mathbb{P}$. Following Serre [21], we call $\mathcal{P}$ regular of density $\alpha>0$ if
@@
By the Chebotarev Density Theorem [21, Proposition 1.5], if $\mathcal{P}$ is Frobenian of density $\alpha>0$ then it is regular of density $\alpha>0$.
@@
Then $N_{S}(X)=\sum_{n\leq X}a_{n}$. We shall make use of a Tauberian theorem due to Delange [23, page 350] to estimate $N_{S}(X)$.
@@
3. Proof of Theorem 6
@@
Suppose $F$ satisfies the hypotheses of Theorem 6. Then there is a set $S$ of primes having positive regular density, such that for all $t\in\mathbb{Q}-\Theta$ and all $p\in S$, the valuation $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))$ is even.
@@
$\Delta(\mathcal{O}_{K_{t}})/\Delta_{Y}(F(t,Y))\;\in\;(\mathbb{Q}^{\times})^{2}.$
@@
$\Delta(\mathcal{O}_{K_{t}})/g(t)\;\in\;(\mathbb{Q}^{\times})^{2}.$
@@
$\Delta(\mathcal{O}_{K_{t}})/G(a,b)\;\in\;(\mathbb{Q}^{\times})^{2}.$
@@
$\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))\;\equiv\;\operatorname{ord}_{p}(G(a,b))\pmod{2}$.
By assumption, there is some $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$. Therefore, by the Chebotarev Density Theorem, there is a set of primes $S$, having positive regular density, such that the congruence $G(u,v)\equiv 0\pmod{p}$ does not have solutions with coprime $u$, $v$. Hence for $p\in S$, we have $\operatorname{ord}_{p}(G(a,b))=0$, so $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))$ is even. ∎
@@
4. Counting low degree fields with local restrictions
In this section we recount some well-known theorems which count number fields of degrees $2$, 33, 44, 55, with certain local restrictions. We shall not need the most general form of these theorems, and content ourselves with stating simplified special cases. The first concerns quadratic fields, and is a special case of [7, Proposition A.1].
Theorem 9 (Ellenberg, Pierce and Wood).
Let $T$ be a finite set of primes. Write $N_{T}^{\pm}(X)$ for the number of isomorphism classes of quadratic fields $K$ satisfying
  • $0<\pm\Delta_{K}\leq X$;
  • all $p\in T$ are unramified in $K$.
Then
$N_{T}^{\pm}(X)\;=\;C_{T}\cdot X+O_{T}(X^{1/2})$
where $C=1/2\zeta(2)$ and
$C_{T}\;=\;C\cdot\prod_{p\in T}\left(1-\frac{1}{p+1}\right).$
Most importantly for us, we shall need a theorem of Bhargava, Taniguchi and Thorne [1, Theorem 1.3] that counts cubic fields, ordered by discriminant, with local specifications. The following two corollaries are easy consequences of that theorem.

### v2

@@ Abstract.
Let $C/\mathbb{Q}$ be a smooth projective absolutely irreducible curve of genus $\geq 2$. For an extension $L/\mathbb{Q}$, let $C(L)_{\mathrm{new}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{P\in C(L)\;:\;\mathbb{Q}(P)=L\}$; this is the set of points on $C$ defined over $L$, but not over any strictly smaller field. Let $n\geq 2$. We conjecture that $C(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of degree $n$ number fields $L$, when ordered by absolute discriminant. For degrees $n=2$, 33, we give sufficient criteria for our conjecture to hold in terms of an explicit model for $C$. For general $n$ we prove a theorem that harmonises with the conjecture. In particular, we verify our conjecture in these cases
  • $n=2$; $C=X_{0}(N)$ for the $18$ values $N\neq 37$ such that $X_{0}(N)$ is hyperelliptic.
  • $n=3$; $C=X_{0}(23)$, $X_{0}(29)$, $X_{0}(31)$, $X_{0}(64)$.
Moreover, we prove the analogue of our conjecture for the unit equation, again with $n=3$.
@@ 1. Introduction
Let $C$ be a smooth projective and absolutely irreducible curve of genus $g$, defined over a number field $K$. Let $D$ be a reduced effective divisor on $C$ and consider the punctured curve X=CDX=C-D. The Euler characteristic of $X$ is given by $\chi(X)=2-2g-\deg(D)$. We say that $X$ is hyperbolic if $\chi(X)<0$. We write $\mathcal{O}_{K}$ for the ring of integers of $K$, and we let $\mathscr{X}/\mathcal{O}_{K}$ be a model for $X$ over $\mathcal{O}_{K}$. The famous Faltings–Siegel Theorem asserts that $\mathscr{X}(\mathcal{O}_{K})$ is finite. We mention three special cases.
@@
  • Let $X=\mathbb{P}^{1}-\{0,1,\infty\}$, which has Euler characteristic 1-1 and is therefore hyperbolic. It can be shown that $\varepsilon\in\mathscr{X}(\mathcal{O}_{K})$ if and only if $\varepsilon$ and $1-\varepsilon$ are units. Thus Faltings–Siegel generalises Siegel’s Theorem which states that the unit equation
@@
A typical instance of the above question is when $X=C$ is a complete curve of genus $g\geq 2$, and $\mathcal{L}$ is the set of all extensions $L/K$ of a fixed degree $n$. This instance of the question is well-considered in the literature; for an excellent survey we recommend [35], but merely mention two famous results.
@@
  • A theorem of Harris and Silverman [15] asserts that if a (complete) curve $C/\mathbb{Q}$ of genus $\geq 2$ has infinitely many quadratic points then $C$ is hyperelliptic or bielliptic. This theorem builds on a deep theorem of Faltings asserting that if $A/K$ is an abelian variety and $V$ is a subvariety not containing any abelian variety then $V(K)$ is finite.
In this paper we offer a conjectural statistical (partial) response to Question 1; this is Conjecture 2 stated below. Let $X/K$ be a hyperbolic curve as before, and let $L/K$ be an extension. The conjecture is concerned with what we call the set of $L$-new points that appear over $L$ on $X/K$; we define these by
@@
We note that the $L$-new points are ones defined over $L$, but not over any strictly smaller extension of $K$. Before stating our conjecture, we explain what we mean by a family $\mathcal{L}$ of extensions $L/K$.
@@
Let $K$ be a number field. Let $n\geq 2$. Let $G_{1},\dotsc,G_{r}$ be transitive subgroups of $S_{n}$. Let $\upsilon_{1},\dotsc,\upsilon_{s}$ be a finite number of places of $K$, and let $L_{i}/K_{\upsilon_{i}}$ be degree $n$ étale algebras. By the degree $n$ family $\mathcal{L}$ with data $(G_{1},\dotsc,G_{r};L_{1},\dotsc,L_{s})$ we mean all extensions $L/K$ such that the following two conditions hold.
@@
If $X=C$ is a smooth projective absolutely irreducible curve over a number field $K$, and $L/K$ is a finite extension, then
@@
  • It is tempting to conjecture that $\mathscr{X}(\mathcal{O}_{L})=\mathscr{X}(\mathcal{O}_{K})$ for $100\%$ of $L\in\mathcal{L}$. This however is false. For example, let $\mathcal{L}$ be all quartic number fields $L/\mathbb{Q}$. Let $X/\mathbb{Q}$ be a hyperbolic curve such that $\mathscr{X}(\mathbb{Z})=\emptyset$, but $\mathscr{X}(\mathbb{Z}[(1+\sqrt{5})/2])\neq\emptyset$. A positive proportion of quartic number fields $L$, when ordered by absolute discriminant, contain $\mathbb{Q}(\sqrt{5})$ ([5, Corollary 1.2]) and therefore $\mathscr{X}(\mathcal{O}_{L})\neq\mathscr{X}(\mathbb{Z})$ for those fields.
@@
The purpose of this paper is to prove Conjecture 2 for several hyperbolic curves, with $\mathcal{L}$ the set of all quadratic or cubic fields.
Theorem 4.
Let $g\in\mathbb{Q}[T]$ be squarefree, of even degree $d\geq 6$, and write $G=\operatorname{Gal}(g)$ for the Galois group of $g$. Suppose there is an element $\sigma\in G$ acting freely on the roots of $g$. Let $C/\mathbb{Q}$ be the hyperelliptic curve defined by $Y^{2}=g(T)$, and write $J$ for its Jacobian. Suppose
  1. (a)
    either $d\geq 10$;
  2. (b)
    or $d=6$ and $J(\mathbb{Q})$ is finite;
  3. (c)
    or $d=8$ and $J(\mathbb{Q})$ is finite or $J$ is simple.
Then $C(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of quadratic fields $L/\mathbb{Q}$, when ordered by absolute discriminant.
We note that if $g$ is irreducible, then there is an element $\sigma\in G$ acting freely on its roots [17]. Naturally, as $C$ is hyperelliptic, it has infinitely many quadratic points, but the above theorem is saying that the quadratic fields over which they appear are relatively rare.
A famous theorem of Ogg [27] asserts that there are $19$ values of $N$ for which which $X_{0}(N)$ is hyperelliptic:
  • genus $2$: $N=22$, $23$, $26$, $28$, $29$, $31$, $37$, $50$;
  • genus $3$: $N=30$, $33$, $35$, $39$, $40$, $41$, $48$;
  • genus $4$: $N=47$;
  • genus $5$: $N=46$, $59$;
  • genus $6$: $N=71$.
Corollary 5.
Let $N\neq 37$ be any of the above values for which $X_{0}(N)$ is hyperelliptic. Then $X_{0}(N)(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of quadratic fields $L/\mathbb{Q}$, when ordered by absolute discriminant.
Theorem 6.
Let $F(T,Y)\in\mathbb{Q}[T,Y]$ have degree $3$ in $Y$, and suppose $F$ is irreducible over $\overline{\mathbb{Q}}$. Let $C/\mathbb{Q}$ be the smooth projective curve with plane affine model $F(T,Y)=0$. Suppose $\operatorname{genus}(C)\geq 8$. Let $\Delta_{Y}(F)\in\mathbb{Q}[T]$ be the discriminant of $F$ with respect to $Y$, and write this as
$\Delta_{Y}(F)\;=\;g(T)\cdot h(T)^{2}$
where $g$, $h\in\mathbb{Q}[T]$ with $g$ squarefree. Suppose that $g$ has even degree, and that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$. Then $C(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of cubic fields $L/\mathbb{Q}$, when ordered by absolute discriminant.
We note that the curve $C$ in Theorem 6 has infinitely many cubic points. Theorems 7 and  8 concern cubic points on certain hyperbolic curves, all of which have infinitely many cubic points.
@@
The four modular curves in Theorem 8 do not satisfy the hypothesis of Theorem 6. Nevertheless, the result follows from the proof of Theorem 6 together with further computations.
Theorems 67 and 8 make use of results due to Bhargava, Taniguchi and Thorne, recalled in Section 5, which count (by absolute discriminant) cubic number fields satisfying local restrictions. For number fields of general degree, there are no such results yet. However, in certain settings we are able to say something about the discriminants of the primitive degree $n$ number fields for which there are new points. Recall that a degree $n$ number field $L/\mathbb{Q}$ is primitive if its Galois group is a primitive subgroup of $S_{n}$; equivalently, $L/\mathbb{Q}$ is primitive if the only subfields $\mathbb{Q}\subseteq K\subseteq L$ are $K=\mathbb{Q}$ and $K=L$.
@@
Let $n\geq 2$. Let $F(T,Y)\in\mathbb{Q}[T,Y]$ have degree $n$ in $Y$, and suppose $F$ is irreducible over $\overline{\mathbb{Q}}$. Let $C/\mathbb{Q}$ be the smooth projective curve with plane affine model $F(T,Y)=0$. Suppose
(3) $\operatorname{genus}(C)\;>\;n^{2}-n+1.$
Let $\Delta_{Y}(F)\in\mathbb{Q}[T]$ be the discriminant of $F$ with respect to $Y$, and write this as
@@
where $g$, $h\in\mathbb{Q}[T]$ with $g$ squarefree. Suppose that $g$ has even degree, and that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$. Let $\mathcal{L}$ be the set of all degree $n$ primitive number fields $L/\mathbb{Q}$. Let
$\operatorname{Disc}(C,\mathcal{L})=\{\lvert\Delta(\mathcal{O}_{L})\rvert\;:\;\text{$L\in\mathcal{L}$ and $C(L)_{\mathrm{new}}\neq\emptyset$}\}.$
Then $\operatorname{Disc}(C,\mathcal{L})$ has density 0 among the natural numbers. More precisely, there is some $0<\alpha<1$ such that
$(\operatorname{Disc}(C,\mathcal{L})\cap[1,X])\;=\;O\left(\frac{X}{(\log{X})^{\alpha}}\right)$
as $X\rightarrow\infty$.
The paper is organized as follows. Section 2 counts, using a theorem of Delange, integers where the prime factors of the squarefree part belongs to certain special subsets of primes. In Section 3 we consider $F(T,Y)\in\mathbb{Q}[T,Y]$ of degree $n$ in $Y$ and, under certain hypotheses, we give an upper bound for the number of discriminants of degree $n$ number fields arising from polynomials $F(t,Y)$ with $t\in\mathbb{Q}$. We apply this in Section 4, together with a theorem of Vojta, to deduce Theorem 9 and Theorem 4. In Section 5 we recount theorems due to Bhargava, Taniguchi and Thorne that count cubic number fields with certain restrictions on ramified primes; these together with techniques introduced earlier in the paper allow us to prove Theorem 6. The proofs of Theorems 7 and  8 are given in Sections 6 and  7 respectively, and make use of results proved in Section 5 together with explicit computations. Finally, in Section 8 we suggest research projects, based on the ideas of this paper, that some readers might be interested to pursue.
@@ 2. Counting integers with certain restricted factorisations
Let $\mathbb{P}$ be the set of prime numbers and let $\mathcal{P}\subseteq\mathbb{P}$. Following Serre [29], we call $\mathcal{P}$ regular of density $\alpha>0$ if, for $\operatorname{Re}(s)>1$,
@@
By the Chebotarev Density Theorem [29, Proposition 1.5], if $\mathcal{P}$ is Frobenian of density $\alpha>0$ then it is regular of density $\alpha>0$.
@@
Then $N_{S}(X)=\sum_{n\leq X}a_{n}$. We shall make use of a Tauberian theorem due to Delange [32, page 350] to estimate $N_{S}(X)$.
@@
3. Degree $n$ points from a single source
To motivate what comes next, we mention the following “single source theorem” due to the authors [19, Theorem 1.1]. In this context, by a primitive degree $n$ point on $C/K$, we mean an algebraic point $P\in C(\overline{K})$ such that $K(P)/K$ is a primitive extension of degree $n$.
Theorem.
This theorem is called the “Single Source Theorem” since, with finitely many exceptions, all primitive degree $n$ points come from a single source which is the morphism $\varphi:C\rightarrow\mathbb{P}^{1}$. A variant of this theorem is due to Derickx [8].
In view of our Conjecture 3, it makes sense to consider the family of degree $n$ number fields we obtain from the fibres of this morphism, and to attempt to show that this family contains $0\%$ of degree $n$ number fields. We shall give a result in this spirit.
Let $C$ be a smooth projective absolutely irreducible curve over $\mathbb{Q}$, and let $\varphi:C\rightarrow\mathbb{P}^{1}$ be a morphism of degree $n\geq 2$. Then the curve $C$ possesses a plane affine (possibly singular) model
C:F(T,Y)=0C\;:\;F(T,Y)=0
where $F(T,Y)\in\mathbb{Q}[T,Y]$ has degree $n$ in $Y$, is irreducible over $\overline{\mathbb{Q}}$, and where the morphism $\varphi$ is given by the map $(T,Y)\mapsto T$. Let $\Theta\subset\mathbb{Q}$ be the set of $t\in\mathbb{Q}$ such that $F(t,Y)$ is either reducible, or has degree <n<n. Hilbert’s irreducibility theorem asserts that $\Theta$ is contained in a thin set. Thus $C$ has infinitely many degree $n$ points. For $t\in\mathbb{Q}-\Theta$ we write $L_{t}$ for the degree $n$ number field defined by $F(t,Y)=0$, and let $\Delta(\mathcal{O}_{L_{t}})$ be the discriminant of the ring of integers $\mathcal{O}_{L_{t}}$ of $L_{t}$. We define the discriminant set of $F$ to be
$\operatorname{Disc}_{F}\;=\;\{\lvert\Delta(\mathcal{O}_{L_{t}})\rvert\;:\;t\in\mathbb{Q}-\Theta\}.$
Theorem 11.
$\Delta_{Y}(F)\;=\;g(T)\cdot h(T)^{2}$
where $g$, $h\in\mathbb{Q}[T]$, with $g$ squarefree. Suppose $g$ has even degree. Let $G=\operatorname{Gal}(g)$. Suppose that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$. Then there is some $0<\alpha<1$ such that, as $X\rightarrow\infty$,
@@ 3. Degree $n$ points from a single source
Suppose $F$ satisfies the hypotheses of Theorem 11. Then there is a set $S$ of primes having positive regular density, such that for all $t\in\mathbb{Q}-\Theta$ and all $p\in S$, the valuation $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{L_{t}}))$ is even.
@@
$\Delta(\mathcal{O}_{L_{t}})/\Delta_{Y}(F(t,Y))\;\in\;(\mathbb{Q}^{\times})^{2}.$
@@
$\Delta(\mathcal{O}_{L_{t}})/g(t)\;\in\;(\mathbb{Q}^{\times})^{2}.$
@@
$\Delta(\mathcal{O}_{L_{t}})/G(a,b)\;\in\;(\mathbb{Q}^{\times})^{2}.$
@@
$\operatorname{ord}_{p}(\Delta(\mathcal{O}_{L_{t}}))\;\equiv\;\operatorname{ord}_{p}(G(a,b))\pmod{2}$.
By assumption, there is some $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$. Therefore, by the Chebotarev Density Theorem, there is a set of primes $S$, having positive regular density, such that the congruence $G(u,v)\equiv 0\pmod{p}$ does not have solutions with coprime $u$, $v$. Hence for $p\in S$, we have $\operatorname{ord}_{p}(G(a,b))=0$, so $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{L_{t}}))$ is even. ∎
@@
4. Proofs of Theorems 4 and 9
We shall need the following theorem of Vojta [36, Corollary 0.3].
Theorem (Vojta).
Let $C$ be a curve defined over a number field $K$. Let $n$ be a positive integer and let $\varphi:C\rightarrow\mathbb{P}^{1}$ be a non-constant morphism defined over $K$. Suppose
(6) $\operatorname{genus}(C)-1\;>\;(n-1)\cdot\deg(\varphi).$
Then the set
$\{P\in C(\overline{K})\;:\;\text{$[K(P):K]\leq n$ and $K(\varphi(P))=K(P)$}\}$
is finite.
Proof of Theorem 9.
Let $C/\mathbb{Q}$ and $F(T,Y)\in\mathbb{Q}[T,Y]$ and $\mathcal{L}$ be as in the statement of Theorem 9. In particular $F$ has degree $n$ in $Y$. Let $\varphi:C\rightarrow\mathbb{P}^{1}$ be defined on the affine patch $F(T,Y)=0$ by $(T,Y)\mapsto T$; this has degree $n$. Let $L\in\mathcal{L}$. Thus $L$ is a primitive degree $n$ number field. Let $P\in C(L)_{\mathrm{new}}$. Then $\mathbb{Q}(P)=L$. Moreover, $\mathbb{Q}(\varphi(P))\subseteq L$, and therefore, as $L$ is primitive, $\mathbb{Q}(\varphi(P))=\mathbb{Q}$ or $L$. We note that assumption (3) in Theorem 9 implies (6). It follows from Vojta’s theorem that for all but finitely many $L\in\mathcal{L}$, any $(T,X)\in C(L)_{\mathrm{new}}$ satisfies $T\in\mathbb{Q}$. Thus, with at most finitely many exceptions, the elements of $\operatorname{Disc}(C,\mathcal{L})$ belongs to $\operatorname{Disc}_{F}$. Theorem 9 follows immediately from Theorem 11. ∎
Proof of Theorem 4.
Let $F(T,Y)=Y^{2}-g(T)$; the discriminant with respect to $Y$ is $\Delta_{Y}(F)=2^{2}g(T)$. Let $d=\deg(g)$. Suppose first that $d\geq 10$. Thus $C$ has genus $\geq 4$, and so (3) is satisfied. Theorem 4 follows immediately from Theorem 9.
We now consider $d=6$ and $d=8$ (and thus $C$ has genus $2$ or $3$). We suppose that assumptions (b), (c) in the statement of Theorem 4 hold. We invoke [18, Proposition 18], which is a standard consequence of the proof by Faltings [10] of the Bombieri-Lang conjecture for subvarieties of abelian varieties. We conclude the existence of finitely many effective degree $2$ divisors $D_{1},\dots,D_{m}$ such that
$C^{(2)}(\mathbb{Q})=\bigcup_{i=1}^{m}\lvert D_{i}\rvert$
where $\lvert D\rvert$ denotes the complete linear series corresponding to $D$. An easy application of the Riemann-Roch theorem and Clifford’s theorem [16, Theorem IV.5.4] shows that $\lvert D\rvert=\{D\}$ unless $D$ is a hyperelliptic divisor, in which case $\lvert D\rvert=\{\varphi^{*}(t):t\in\mathbb{P}^{1}(\mathbb{Q})\}$. Thus
$\{P\in C(\overline{\mathbb{Q}})\;:\;\text{$[\mathbb{Q}(P):\mathbb{Q}]=2$ and $\mathbb{Q}(\varphi(P))=\mathbb{Q}(P)$}\}$
is finite. The theorem now follows from the proof of Theorem 9. ∎
Proof of Corollary 5.
It is known, e.g. [4], that $J_{0}(N)(\mathbb{Q})$ is finite for all NN such that $X_{0}(N)$ is hyperellitic, except for $N=37$. The hyperelliptic polynomials $g$ for these hyperelliptic modular curves are available in the computer algebra package Magma [2]. We computed the Galois groups and checked that the hypotheses of Theorem 4 apply. ∎
Remark. A recent paper of Najman and Trbović [25] studies primes that ramify in the field of definition of non-exceptional non-cuspidal quadratic points on hyperelliptic modular curves $X_{0}(N)$ with $N\neq 37$. It is possible to deduce Corollary 5 from their results, together with Theorem 10.
5. Proof of Theorem 6
We shall need a theorem of Bhargava, Taniguchi and Thorne [1, Theorem 1.3] that counts cubic fields, ordered by discriminant, with local specifications. The following two corollaries are easy consequences of that theorem.